OpenStudy (anonymous):

What is 3 √250 expressed in simplest radical form?

OpenStudy (jdoe0001):

 thus \(\bf 250 = 2 \times 5 \times 5 \times 5 \implies 2 \times 5^2 \times 5\\\quad \\ 3\sqrt{250} \implies 3\sqrt{2 \times 5^2 \times 5}\)